OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) the sum of a number and its reciprocal is 73/24. Find the number..

OpenStudy (anonymous):

x+(1/x)=73/24 x^2+1=(73/24)x x^2-(73/24)x+1=0 now solve for the quadratic and you get x=8/3, x=3/8